a large kitchen pot initially contains 7.5 kg of water at a burner is turned on and heated

heat Is 1.2.6. A large kitchen pot initially contains 7.5 kg of water at 20°C. A burner is turned on and added to the water at a rate of 500 cal/s. Determine how many minutes it will take for one-third of the mass to evaporate at standard atmospheric pressure.
Answer #1

Specific heat of water C = 4186 K / kgC

Latent heat of vaporization of water L = 2260 * 103 J/kg

Amount of heat required to raise temperature of m - 7.5 kg of water at 20° to 100° and then evaporating one-third 2.5kg) of water is Q = mCAT + ma evaporated

Q = 7.5 * 4186 * 80+ 2.5 * 2260 * 103-8161600 J

Rate at which heat is added is dQ 500 cal s 2092 Js lt

Time taken for the mass to evaporate to one-third is Q8161600 3901 s-65 min ag-2092